OpenStudy (anonymous):

Solve for x and y using elimination. −4x−2y=4 −6x−7y=2

OpenStudy (anonymous):

first equation x3,second equation x2?

OpenStudy (campbell_st):

the 1st equation can be written as -2x - y = 2 call this (1) -6x - 7y = 2 (2) multiply equation (1) by -3 and add to equation 2 6x +3y = -6 -6x - 7y = 2 ______________ -4y = -4 then y = 1 sub into equation (1) to find x